# Concepts in Motion: Temporal Bottlenecks for Interpretable Video Classification

### v1

@@ 3 Experiments
In this section, we evaluate MoTIF on known video classification benchmarks. We then analyze different architectural design choices through ablation studies, illustrate the kinds of explanations our model provides with representative examples, and finally examine its behavior under interventions.
@@ 3.1 Experimental Setup
Backbones. We adopt a range of CLIP-based backbones differing in model size, patch resolution, and temporal adaptation as feature extractors. Specifically, we include RN/50, ViT-B/16, ViT-B/32, and ViT-L/14 from CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), and video-adapted Perception Encoder (PE) (Bolya et al., 2025). SigLIP replaces CLIP’s contrastive softmax with a pairwise sigmoid loss, decoupling image and text encoders, which improves training efficiency and scalability. The Perception Encoder is trained on video–text pairs and incorporates a mechanism to aggregate frame-level information. For both PE and SigLIP we use the ViT-L/14 backbone.
Concepts. For concept creation, we follow Yang et al. (2023) and prompt an LLM (GPT-5) (OpenAI et al., 2024) to generate unique textual concepts per dataset; these concepts are further described in Appendix F.
Baselines. In line with (Rao et al., 2024; Prasse et al., 2025), we evaluate MoTIF against zero-shot and linear-probe baselines. To ensure a fair comparison with MoTIF, we compute zero-shot predictions on the window-level and apply majority voting across windows, preventing the zero-shot baseline from relying solely on a single global aggregation. As a supervised baseline (linear probe), we train a linear classifier on top of mean-pooled window embeddings, analogous to a global bottleneck. This setup captures only coarse video-level information, in contrast to MoTIF, which models temporally localized concept activations. All experiments were conducted on a single NVIDIA A6000 GPU with 48 GB of VRAM.
@@ 3.2.1 Performance Evaluation
We report Top-1 accuracies of MoTIF with different backbones across four datasets in Table 1, using the uniform hyperparameters described in A.1. For each dataset, we provide the mean and standard deviation over the available test splits (see Appendix D.2). The linear probe consistently outperforms zero-shot, while MoTIF surpasses both. Accuracy increases with backbone capacity, indicating that introducing a concept bottleneck in video models preserves—and can improve—predictive performance while adding interpretability.
Moreover, variants such as SigLIP and PE achieve higher accuracy than their counterparts with comparable parameter counts. Performance is strong across datasets; SSv2 remains the most challenging due to its abstract classes (e.g., putting something onto something), which require MoTIF to capture similar relational representations. On Breakfast, MoTIF outperforms the linear probe by 661616 percentage points; however, all methods (zero-shot, linear probe, and MoTIF) exhibit large standard deviations across test splits. For HMDB51, UCF101, and SSv2, the gap is smaller (0.90.98.28.2 percentage points), likely because these datasets consist of shorter clips, where a simple mean representation already suffices to identify the class, especially for the PE backbone, which shows the strongest performance for video embeddings. However, on some datasets there is still a performance gap compared to strong non-interpretable baselines such as NoFrameLeftBehind (Liu et al., 2021) and VideoMAE V2 (Wang et al., 2023), especially for SSv2 (see Appendix D.1).
Table 1: Performance comparison (% Top-1 accuracy). Mean $\pm$ standard deviation on Breakfast Actions, HMDB51, UCF101, and SSv2 with different CLIP-based backbones.
Method Breakfast HMDB51 UCF101 SSv2
Zero-shot
CLIP-ViT-L/14 (Radford et al., 2021) 31.1 $\pm$ 4.7 45.7 $\pm$ 0.1 70.6 $\pm$ 0.5 2.2
PE-L/14 (Bolya et al., 2025) 41.4 $\pm$ 7.0 56.7 $\pm$ 0.6 74.6 $\pm$ 0.9 0.7
Linear probe
PE-L/14 (Bolya et al., 2025) 72.9 $\pm$ 10.3 74.4 $\pm$ 0.6 94.5 $\pm$ 0.6 25.5
MoTIF (ours)
MoTIF (RN/50) 52.8 $\pm$ 6.9 62.8 $\pm$ 1.1 82.8 $\pm$ 0.6 16.0
MoTIF (ViT-B/32) 53.4 $\pm$ 6.9 65.3 $\pm$ 1.8 85.6 $\pm$ 1.2 17.5
MoTIF (ViT-L/14) 69.3 $\pm$ 6.2 73.3 $\pm$ 1.0 93.2 $\pm$ 0.7 20.4
MoTIF (SigLIP-L/14) 73.5 $\pm$ 8.6 73.2 $\pm$ 2.4 94.0 $\pm$ 0.8 22.4
MoTIF (PE-L/14) 83.6 $\pm$ 6.5 79.6 $\pm$ 0.3 95.4 $\pm$ 0.7 30.0
@@ 3.2.2 Ablations
Figure 3: Effect of log-sum-exp temperature $\tau$ on accuracy and entropy. Accuracy is stable across $\tau$, while both concept- and logit-level entropy decrease as $\tau$ increases, yielding sharper time-importance distributions.
Temperature $\tau$.
We vary the log-sum-exp pooling temperature $\tau$ to assess its effect on both accuracy and the sharpness of the time-importance distribution. Sharpness is quantified by the entropy of the softmax weights, computed either at the concept or at the logits level. Experiments show (see Figure 3) that accuracy varies only slightly across all tested values of $\tau$, indicating robustness of the predictive performance. In contrast, entropy decreases monotonically with larger $\tau$, i.e., the temporal weighting becomes more concentrated on fewer decisive frames. Hence, $\tau$ provides a controllable parameter: small values result in diffuse, high-entropy explanations, whereas large values produce sharp, low-entropy attributions, while accuracy remains relatively stable. Thus this parameter can be tuned by the user.
@@
Attention variant (full vs. diagonal).
While MoTIF enforces concept isolation, we also evaluate full multi-head attention; Figure 4 shows train and test accuracy for both. This test is designed to investigate if there are performance differences between those two versions. Comparable test accuracies with and without diagonal attention show MoTIF preserves generalization despite the interpretability constraint; only training accuracy differs noticeably. We further illustrate the effect on explanations in B.2. When channels are mixed, the most important dimensions appear essentially random, severely reducing interpretability.
Figure 5: Concept set influence. Distribution of test accuracy with five different concept sets.
Concept set influence. To investigate the influence of the concept proposal set on the performance of MoTIF, we prompt GPT-5 five times, each time requesting a distinct set of candidate concepts (with some natural overlap across runs). All resulting concept sets for the five datasets are listed in Appendix F. Our experiments demonstrate that given the same prompt to the LLM, concept set affects test accuracy only moderately, with consistent trends across datasets and backbones (Fig. 5). In line with prior work on concept bottleneck models, more dataset-specific concepts tend to increase accuracy (Rao et al., 2024; Prasse et al., 2025; Schrodi et al., 2025). However, our main motivation here is not to optimize the quality of the concept proposals, but rather to demonstrate that our framework works broadly and robustly across datasets, backbones, and varying concept sets. In addition, we note that kk—the number of concepts retained in the bottleneck—influences the achievable accuracy: very small $k$ restricts expressive power, while very large $k$ may introduce noise and redundancy if concepts are too similar or irrelevant.
@@ 3.2.3 Explanations
Figure 6: MoTIF explanations. Example videos from Breakfast and SSv2 with correct classifications, illustrating the three explanation modes supported by MoTIF (ViT-L14).
A central goal of our approach is to provide interpretable insights into the decision-making process of video classification models. We present further examples for the three explanation modes introduced in Section 2.2: global concept importance, local concepts in decisive windows, and temporal dependencies between concepts. In Figure 6, the first example from Breakfast Action shows preparing a sandwich. Global and temporal concepts highlight bread and bagel. The attention map for bagel reveals strong vertical stripes, meaning many query frames attend to the same key frames—anchor moments where the concept is most pronounced. Thus, bagel is localized in time but influences the entire sequence. The second example, from UCF101, depicts kayaking. Here, the most salient concepts are paddle and kayak, which remain stable over the short clip. The attention map for paddle shows a more uniform distribution across time, with slight emphasis at $u=3$ and $u=6$, indicating consistent expression of the concept rather than isolated peaks. Additional video examples are provided in the supplementary material (Appendix B.1 discusses misclassifications, and Appendix B.2 without diagonal attention), and all data will be released upon acceptance.
3.2.4 Interventions – Concept Removal
To illustrate the effect of concept removal, we revisit Figure 1. Ablating the most influential concept bow by zeroing its activations changes the prediction from the correct class to run (logit 8.20→6.79). Removing windows 1–4, where the bow is handled, shifts the output to talk (logit 6.75).
Table 2: Concept removal. Accuracy after ablating top-$k$ influential concept channels (RN/50).
Dataset $k=0$ $k=1$ $k=2$ $k=3$ $k=4$
Breakfast 1.000 0.514 0.218 0.085 0.039
HMDB51 1.000 0.314 0.125 0.042 0.012
On the dataset level, Table 2 shows that accuracy declines steeply when the top-$k$ channels are removed (zeroed out). This indicates that MoTIF’s decisions rely strongly on a small set of influential concepts. Note that $k=0$ corresponds to perfect accuracy, since we evaluate relative to MoTIF’s own predictions.
@@ 4 Related Work
MoTIF positions itself at the intersection of three research areas: concept bottleneck models, temporal modeling, and video classification with vision–language backbones. At its core, MoTIF belongs to the family of concept bottleneck models (CBMs), which enforce intermediate, interpretable representations. At the same time, it incorporates transformer-based sequence modeling, situating it within the broader line of work on video classification and activity recognition. To the best of our knowledge, no prior work has combined these fields into a unified framework with three complementary views. Below, we review related work and outline how MoTIF extends these directions.
Concept bottleneck models.
@@
CBMs (Koh et al., 2020) predict concepts as intermediate features before the final class prediction (Havasi et al., 2022; Chauhan et al., 2023; Prasse et al., 2025; Yang et al., 2023; Sawada & Nakamura, 2022). Recent research has focused on automatic concept discovery: DCLIP aligns data with CLIP’s vision–language space (Menon & Vondrick, 2023), LaBo queries large language models for diverse candidate concepts (Yang et al., 2023), and DCBM leverages segmentation foundation models to extract object- and part-level concepts (Prasse et al., 2025). Other works have extended this idea to different modalities. For instance, Ismail et al. (2025) adapt the bottleneck principle to protein design, where interpretable biochemical features form the concept space. Wu et al. (2022) extend CBMs to time series data in the medical domain, demonstrating that clinically meaningful temporal features can act as concepts. More recently, Sun et al. (2025) apply the CBM framework to language models, where intermediate concepts correspond to interpretable linguistic or semantic units.
@@
While CBMs have been widely explored in the image domain, their extension to sequential data has seen little attention. Jeyakumar et al. (2022) propose extracting concepts from video descriptions and using them as inputs to a CBM for video classification, but their approach operates on a global level and thus produces concept activations that are not tied to specific parts of a video. In contrast, MoTIF generalizes the bottleneck principle from static images to temporal sequences, enabling reasoning over evolving concept activations on established video benchmarks.
Video classification and action recognition.
@@
Video classification not only assigns a label to an entire sequence but, in the case of action recognition, also requires identifying the actions occurring within it (Pareek & Thakkar, 2021). Both tasks have progressed from CNN-based architectures (Tran et al., 2015; Lin et al., 2019; Liu et al., 2021) to transformer-based (Wang et al., 2024a) models such as TimeSformer (Bertasius et al., 2021) and VideoMAE (Tong et al., 2022), which capture long-range dependencies and leverage large-scale self-supervised pre-training. MoTIF integrates these approaches by adapting a transformer block for temporal modeling while incorporating a 1D-CNN to enable per-channel computations.
Temporal concept modeling.
@@
While concept-based explanations have been widely studied in the image domain (Knab et al., 2025a; Ghorbani et al., 2019), only a few works explore temporal dynamics (Gulshad et al., 2023; Kowal et al., 2024). Ji et al. (2023) propose a spatio-temporal concept framework to analyze representations in 3D ConvNets. PCBEAR (Lee et al., 2025) introduces static and dynamic pose concepts for action recognition, and Saha et al. (2024) extend TCAV to videos by computing concept importance scores across sequences. 3D-ACE explains video models post hoc using user-defined concepts or supervoxels (Ghorbani et al., 2019). In contrast, MoTIF integrates concept-specific temporal attention directly into the predictive model, enabling reasoning over evolving concepts rather than treating them as fixed inputs or external explanations.

### v2

@@ 3 Experiments
In this section, we first evaluate MoTIF on common video classification benchmarks. We then analyze different architectural design choices through ablation studies, illustrate the kinds of explanations our model provides with representative examples, and finally examine its behavior under interventions.
@@ 3.1 Experimental Setup
Backbones. We adopt a range of CLIP-based backbones differing in model size, patch resolution, and temporal adaptation as feature extractors. Specifically, we include RN/50, ViT-B/32, and ViT-L/14 from CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), and video-adapted Perception Encoder (PE) (Bolya et al., 2025). SigLIP replaces CLIP’s contrastive softmax with a pairwise sigmoid loss, decoupling image and text encoders, which improves training efficiency and scalability. The Perception Encoder is trained on video–text pairs and incorporates a mechanism to aggregate frame-level information. For SigLIP, we use ViT-L/14; for PE, we use both ViT-L/14 and ViT-G/14.
Concept Discovery. For agentic concept discovery (Section 2.1), we use Qwen-3 30B (Yang et al., 2025) to generate the concept sets employed in our main experiments in Table 1. For each dataset, up to five training videos per class are provided to the agent, and concept extraction is performed exclusively on the training split to avoid information leakage. For ablation studies, we follow Yang et al. (2023) and construct smaller, dataset-specific concept sets by prompting an LLM (GPT-5) (OpenAI et al., 2024). Details on the resulting concept vocabularies are provided in Appendix E.
Baselines. In line with (Rao et al., 2024; Prasse et al., 2025), we evaluate MoTIF against zero-shot and linear-probe baselines. To ensure a fair comparison with MoTIF, we compute zero-shot predictions on the window-level and apply majority voting across windows, preventing the zero-shot baseline from relying solely on a single global aggregation. As a supervised baseline (Global CBM), we train a linear classifier on top of mean-pooled window embeddings, analogous to a global bottleneck. This setup captures only coarse video-level information, in contrast to MoTIF, which models temporally localized concept activations. All experiments were conducted on a single H100 GPU with 94 GB of VRAM.
@@ 3.2.1 Performance Evaluation
We report Top-1 accuracies with different backbones across four datasets in Table 1, using the tuned hyperparameters described in Appendix A.1 and agentic concept discovery. The global CBM consistently outperforms zero-shot, while MoTIF surpasses both. Accuracy increases with backbone capacity, suggesting that stronger vision–language representations yield more informative (and thus more easily separable) concept activations, which improves the downstream linear classifier. This indicates that introducing a concept bottleneck can preserve—and sometimes improve—predictive performance while adding interpretability.
Regarding the transformer design, both variants perform well; however, the space–time variant (MoTIF-ST) (Bertasius et al., 2021) consistently achieves higher accuracy on SSv2. Across comparable model sizes, PE-based backbones outperform CLIP variants. Overall performance is strong across datasets: On Breakfast, MoTIF improves over the global CBM by 11111818 percentage points. For HMDB51, UCF101, and SSv2, gains are smaller (0.90.98.28.2 points), likely because these datasets contain short clips where mean pooling is often sufficient—particularly for PE embeddings, which yield the strongest video representations. A performance gap between interpretable concept bottleneck models and strong non-interpretable black-box baselines is primarily observed for VideoMAE V2 (Wang et al., 2023); for TSM (Lin et al., 2019) and NoFrameLeftBehind (Liu et al., 2021), results are similar or better except on SSv2. SSv2 isolates fine-grained relational reasoning between multiple concepts, a setting where enforcing strict concept independence intentionally trades expressivity for interpretability. Despite its explicit focus on interpretability, MoTIF substantially narrows the accuracy gap to black-box models. This is achieved by combining temporally grounded concept modeling with agentic concept discovery, which yields more complete and task-relevant concept sets; an overview of the LLM-based concept set (cf. Table 9) and the corresponding test-split results are provided in Appendix C.1.
@@ 3.2.2 Ablations
Figure 3: Effect of log-sum-exp temperature $\tau$ on accuracy and entropy. Accuracy remains stable, while concept- and logit-level entropy decrease as $\tau$ increases.
Temperature $\tau$. We vary the log-sum-exp pooling temperature $\tau$ to assess its effect on both accuracy and the sharpness of the time-importance distribution. Sharpness is quantified by the entropy of the softmax weights, computed either at the concept or at the logits level. Experiments show (see Figure 3) that accuracy varies only slightly across all tested values of $\tau$, indicating robustness of the predictive performance. In contrast, entropy decreases monotonically with larger $\tau$, i.e., the temporal weighting becomes more concentrated on fewer decisive frames. Hence, $\tau$ provides a controllable parameter: small values result in diffuse, high-entropy explanations, whereas large values produce sharp, low-entropy attributions, while accuracy remains relatively stable. Thus this parameter can be tuned by the user.
@@
Attention variant (full vs. diagonal). While MoTIF enforces concept isolation, we also evaluate a variant with full multi-head attention. Figure 4 compares train and test accuracy for both settings. This comparison explicitly evaluates the impact of cross-concept interactions on both predictive performance and interpretability. Diagonal attention exposes a clear accuracy–interpretability trade-off: enabling cross-concept attention recovers accuracy on demanding datasets, while diagonal attention preserves faithful concept attribution, allowing controlled analysis of how temporal interactions affect both prediction and explanation. On temporally demanding datasets such as SSv2, this trade-off becomes more pronounced, with full attention yielding up to 10.1% higher test accuracy. We further illustrate the effect on explanations in Figure B.2: when channels are mixed, the most influential dimensions appear effectively random, substantially reducing interpretability.
Temporal sensitivity and dynamic concepts. A general challenge in video concept learning is preventing models from defaulting to temporally order-invariant reasoning (Bertasius et al., 2021). To explicitly assess whether MoTIF captures temporal structure, we evaluate temporal dependence in two complementary ways. First, we construct a controlled synthetic benchmark (1,989 sequences, matching the number of Breakfast training instances) in which class identity depends exclusively on frame ordering (ascending, descending, U-shaped, inverted-U, and periodic patterns), such that static appearance cues are uninformative by design (for details see Appendix B.4). Performance collapses from 86.97% on ordered sequences to 21.06% under temporal shuffling (chance $\approx 20\%$), demonstrating that MoTIF explicitly encodes temporal order rather than relying on static appearance. Moreover, on this benchmark the static image-level bottleneck baseline (Global CBM), achieves only 35.5% accuracy, underscoring that temporal modeling is essential when order defines the class.
Table 2: Temporal sensitivity analysis and bottleneck comparison (PE-L/14). Shuffling indicates random permutation of windows at evaluation time; synthetic uses five temporal classes.
Setting Basic Shuffled
Synthetic (MoTIF) 86.97 21.06
Synthetic (Global CBM) 35.5 35.5
Breakfast (MoTIF) 87.3 85.2–86.6
HMDB51 (MoTIF) 79.9 77.1–78.5
UCF101 (MoTIF) 94.7 94.5–94.6
SSv2 (MoTIF) 30.0 26.9–27.4
Second, in Table 2, we apply the same shuffling intervention to real-world datasets. Overall, the degradation is mild across all benchmarks (typically a few percentage points), suggesting that many clips still contain strong appearance cues. SSv2 shows the largest drop (about 2–3 points), which is consistent with its higher reliance on temporal relations; correspondingly, it benefits the most from agent-generated temporal concepts.
@@ 3.2.3 Explanations
Figure 5: MoTIF explanations. Example videos from Breakfast and UCF101 with correct classifications, illustrating the three explanation modes supported by MoTIF (ViT-L14).
A central goal of our approach is to provide interpretable insights into the decision-making process of video classification models. We present further examples for the three explanation modes introduced in Section 2.3: global concept importance, local concepts in decisive windows, and temporal dependencies between concepts. In Figure 5, the first example from Breakfast Action shows preparing a sandwich. Global and temporal concepts highlight bread and bagel. The attention map for bagel reveals strong vertical stripes, meaning many query frames attend to the same key frames—anchor moments where the concept is most pronounced. Thus, bagel is localized in time but influences the entire sequence. The second example, from UCF101, depicts kayaking. Here, the most salient concepts are paddle and kayak, which remain stable over the short clip. The attention map for paddle shows a more uniform distribution across time, with slight emphasis at $u=3$ and $u=6$, indicating consistent expression of the concept rather than isolated peaks. Additional video examples are provided in the supplementary material (Appendix B.1 discusses misclassifications, and B.2 shows examples without diagonal attention); all data will be released upon acceptance.
3.2.4 Concept Interventions
Table 3: Concept interventions. Accuracy after removing or perturbing $k$ concept channels (RN/50).
Dataset $k$ Top-$k$ Removal Random Removal Random Insertion
Breakfast 0 1.000 1.000 1.000
1 0.496 0.989 0.979
2 0.229 0.972 0.947
3 0.085 0.951 0.951
4 0.028 0.923 0.908
HMDB51 0 1.000 1.000 1.000
1 0.603 0.974 0.978
2 0.374 0.963 0.961
3 0.238 0.949 0.945
4 0.142 0.935 0.921
To illustrate the effect of concept manipulation, we revisit Figure 2. Ablating the most influential concept bow by zeroing its activations changes the prediction from the correct class to run (logit 8.20→6.79). Removing windows 1–4, where the bow is handled, shifts the output to talk (logit 6.75). In addition, we evaluate three interventions: (i) top-$k$ concept removal, (ii) random concept removal, and (iii) random insertion of noise into concept activations ($\epsilon\sim\mathcal{N}(0,\,\sigma^{2})$ with $\sigma=0.5$). On the dataset level, Table 3 shows that accuracy declines sharply only when the most influential concepts are removed, while random removal or noise insertion cause minimal degradation. This indicates that MoTIF’s predictions rely on a small, semantically meaningful subset of concepts and remain robust under noise [^2] Note that $k=0$ corresponds to perfect accuracy, since all values are normalized relative to MoTIF’s baseline predictions..
@@ 4 Related Work
MoTIF lies at the intersection of three research areas: concept bottleneck models, temporal modeling, and video classification with vision–language backbones. At its core, MoTIF belongs to the family of concept bottleneck models (CBMs), which enforce intermediate, interpretable representations, while incorporating transformer-based sequence modeling that situates it within video classification and activity recognition. Below, we review related work and outline how MoTIF extends these directions.
@@
Concept bottleneck models. CBMs (Koh et al., 2020) predict concepts as intermediate features before the final class prediction (Havasi et al., 2022; Chauhan et al., 2023; Prasse et al., 2025; Yang et al., 2023; Sawada and Nakamura, 2022). Recent research has focused on automatic concept discovery: DCLIP aligns data with CLIP’s vision–language space (Menon and Vondrick, 2023), LaBo queries large language models for diverse candidate concepts (Yang et al., 2023), and DCBM leverages segmentation foundation models to extract object- and part-level concepts (Prasse et al., 2025). Other works have extended this idea to different modalities. For instance, Ismail et al. (2025) adapt the bottleneck principle to protein design, where interpretable biochemical features form the concept space. Wu et al. (2022) extend CBMs to medical time series data, demonstrating that clinically meaningful temporal features can act as concepts. More recently, Sun et al. (2025) apply the CBM framework to language models, where intermediate concepts correspond to interpretable linguistic or semantic units.
@@
While CBMs have been widely explored in the image domain, their extension to sequential data has seen little attention. Jeyakumar et al. (2022) propose extracting concepts from video descriptions and using them as inputs to a CBM for video classification, but their approach operates on a global level and thus produces concept activations that are not tied to specific parts of a video. Lee et al. (2025a) propose DANCE, which disentangles video concepts by combining pose sequences with textual concepts, providing users with a two-sided comparison between textual concepts and motion-dynamic concepts. In contrast, MoTIF generalizes the CBM principle from static images to arbitrary-length temporal sequences and focuses on agentic discovery of purely semantic textual concepts grounded in video data.
@@
Video classification and action recognition. Video classification not only assigns a label to an entire sequence but, in the case of action recognition, also requires identifying the actions occurring within it (Pareek and Thakkar, 2021). Both tasks have progressed from CNN-based architectures (Tran et al., 2015; Lin et al., 2019; Liu et al., 2021) to transformer-based (Wang et al., 2024a) models such as TimeSformer (Bertasius et al., 2021) and VideoMAE (Tong et al., 2022), which capture long-range dependencies and leverage large-scale self-supervised pre-training. MoTIF integrates these approaches by adapting a transformer block for temporal modeling while incorporating a 1D-CNN to enable per-channel computations.
@@
Temporal concept modeling. While concept-based explanations have been widely studied in the image domain (Knab et al., 2025a; Ghorbani et al., 2019), only a few works explore temporal dynamics (Gulshad et al., 2023; Kowal et al., 2024). Ji et al. (2023) propose a spatio-temporal concept framework to analyze representations in 3D ConvNets. PCBEAR (Lee et al., 2025b) introduces static and dynamic pose concepts for action recognition, and Saha et al. (2024) extend TCAV to videos by computing concept importance scores across sequences. 3D-ACE explains video models post hoc using user-defined concepts or supervoxels (Ghorbani et al., 2019). In contrast, MoTIF integrates concept-specific temporal attention directly into the predictive model, enabling reasoning over evolving concepts rather than treating them as fixed inputs or external explanations.